\begin{thebibliography}{17}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Felice(1971)]{Felice1971}
F.~Felice.
\newblock {On the gravitational field acting as an optical medium}.
\newblock \emph{General Relativity and Gravitation}, 2\penalty0 (4):\penalty0
  347--357, 1971.
\newblock ISSN 0001-7701.
\newblock \doi{10.1007/BF00758153}.
\newblock URL \url{http://link.springer.com/10.1007/BF00758153}.

\bibitem[Korol’kova et~al.(2013)Korol’kova, Kulyabov, and
  Sevast’yanov]{kulyabov:2013:springer:cadabra}
A.~V. Korol’kova, D.~S. Kulyabov, and L.~A. Sevast’yanov.
\newblock {Tensor computations in computer algebra systems}.
\newblock \emph{Programming and Computer Software}, 39\penalty0 (3):\penalty0
  135--142, 2013.
\newblock ISSN 0361-7688.
\newblock \doi{10.1134/S0361768813030031}.
\newblock URL \url{http://link.springer.com/10.1134/S0361768813030031}.

\bibitem[Kulyabov(2013)]{kulyabov:2013:conf:maxwell}
D.~S. Kulyabov.
\newblock {Geometrization of Electromagnetic Waves}.
\newblock In \emph{Mathematical Modeling and Computational Physics}, page 120,
  Dubna, 2013. JINR.
\newblock ISBN 978-5-9530-0362-9.
\newblock URL \url{http://mmcp2013.jinr.ru}.

\bibitem[Kulyabov and
  Nemchaninova(2011)]{kulyabov:2011:vestnik:curve-maxwell::en}
D.~S. Kulyabov and N.~A. Nemchaninova.
\newblock Maxwell’s equations in curvilinear coordinates.
\newblock \emph{Bulletin of Peoples’ Friendship University of Russia. Series
  Mathematics. Information Sciences. Physics}, \penalty0 (2):\penalty0
  172--179, 2011.
\newblock in Russian.

\bibitem[Kulyabov et~al.(2012)Kulyabov, Korolkova, and
  Korolkov]{kulyabov:2012:vestnik:2012-1}
D.~S. Kulyabov, A.~V. Korolkova, and V.~I. Korolkov.
\newblock {Maxwell’s Equations in Arbitrary Coordinate System}.
\newblock \emph{Bulletin of Peoples’ Friendship University of Russia. Series
  ``Mathematics. Information Sciences. Physics''}, \penalty0 (1):\penalty0
  96--106, 2012.
\newblock URL \url{http://arxiv.org/abs/1211.6590}.

\bibitem[Leonhardt and Philbin(2009)]{leonhardt:2009:light}
U.~Leonhardt and T.~G. Philbin.
\newblock {Transformation optics and the geometry of light}.
\newblock In \emph{Progress in Optics}, volume~53, pages 69--152. 2009.
\newblock \doi{10.1016/S0079-6638(08)00202-3}.
\newblock URL
  \url{http://www.sciencedirect.com/science/article/pii/S0079663808002023}.

\bibitem[Leonhardt et~al.(2008)Leonhardt, Philbin, and Haugh]{Leonhardt2008}
U.~Leonhardt, T.~G. Philbin, and N.~Haugh.
\newblock {General Relativity in Electrical Engineering}.
\newblock pages 1--19, 2008.

\bibitem[Minkowski(1908)]{minkowski:1908}
H.~Minkowski.
\newblock {Die Grundlagen f\"{u}r die electromagnetischen Vorg\"{a}nge in
  bewegten K\"{o}rpern}.
\newblock \emph{Nachrichten von der Gesellschaft der Wissenschaften zu
  G\"{o}ttingen, Mathematisch-Physikalische Klasse}, \penalty0 (68):\penalty0
  53--111, 1908.

\bibitem[Penrose and Rindler(1984)]{penrose-rindler-1987::en}
R.~Penrose and W.~Rindler.
\newblock \emph{{Spinors and Space-Time: Two-Spinor Calculus and Relativistic
  Fields}}, volume~1.
\newblock Cambridge University Press, 1984.

\bibitem[Plebanski(1960)]{Plebanski1960}
J.~Plebanski.
\newblock {Electromagnetic waves in gravitational fields}.
\newblock \emph{Physical Review}, 118\penalty0 (5):\penalty0 1396--1408, 1960.
\newblock \doi{10.1103/PhysRev.118.1396}.
\newblock URL \url{http://prola.aps.org/abstract/PR/v118/i5/p1396_1}.

\bibitem[Sivukhin(1979)]{sivukhin:1979:ufn::en}
D.~V. Sivukhin.
\newblock {The international system of physical units}.
\newblock \emph{Soviet Physics Uspekhi}, 22\penalty0 (10):\penalty0 834--836,
  1979.
\newblock \doi{10.1070/PU1979v022n10ABEH005711}.
\newblock URL \url{http://ufn.ru/en/articles/1979/10/g/}.

\bibitem[Sommerfeld(1964)]{sommerfeld1964lectures}
A.~Sommerfeld.
\newblock \emph{{Lectures on Theoretical Physics: Electrodynamics}}.
\newblock Lectures on Theoretical Physics. Academic Press, 1964.
\newblock URL \url{http://books.google.ru/books?id=wMnvAAAAMAAJ}.

\bibitem[Stratton(1941)]{stratton:1948::en}
J.~A. Stratton.
\newblock \emph{{Electromagnetic Theory}}.
\newblock MGH, 1941.

\bibitem[Tamm(1924)]{tamm:1924:jrpc::en}
I.~E. Tamm.
\newblock {Electrodynamics of an anisotropic medium in a special theory of
  relativity}.
\newblock \emph{Russian Journal of Physical and Chemical Society. Part
  physical}, 56\penalty0 (2-3):\penalty0 248--262, 1924.
\newblock in Russian.

\bibitem[Tamm(1925)]{tamm:1925:jrpc::en}
I.~E. Tamm.
\newblock {Crystal optics theory of relativity in connection with geometry
  biquadratic forms}.
\newblock \emph{Russian Journal of Physical and Chemical Society. Part
  physical}, 57\penalty0 (3-4):\penalty0 209--240, 1925.
\newblock in Russian.

\bibitem[Tamm and Mandelstam(1925)]{tamm:1925:mathann}
I.~E. Tamm and L.~I. Mandelstam.
\newblock {Elektrodynamik der anisotropen Medien in der speziellen
  Relativitatstheorie}.
\newblock \emph{Mathematische Annalen}, 95\penalty0 (1):\penalty0 154--160,
  1925.

\bibitem[Thompson et~al.(2011)Thompson, Cummer, and Frauendiener]{Thompson2011}
R.~T. Thompson, S.~A. Cummer, and J.~Frauendiener.
\newblock {A completely covariant approach to transformation optics}.
\newblock \emph{Journal of Optics}, 13\penalty0 (2):\penalty0 024008, 2011.
\newblock ISSN 2040-8978.
\newblock \doi{10.1088/2040-8978/13/2/024008}.

\end{thebibliography}


\begin{thebibliography}{10}
\def\selectlanguageifdefined#1{
\expandafter\ifx\csname date#1\endcsname\relax
\else\selectlanguage{#1}\fi}
\providecommand*{\href}[2]{{\small #2}}
\providecommand*{\url}[1]{{\small #1}}
\providecommand*{\BibUrl}[1]{\url{#1}}
\providecommand{\BibAnnote}[1]{}
\providecommand*{\BibEmph}[1]{#1}
\ProvideTextCommandDefault{\cyrdash}{\hbox to.8em{--\hss--}}
\providecommand*{\BibDash}{\ifdim\lastskip>0pt\unskip\nobreak\hskip.2em\fi
\cyrdash\hskip.2em\ignorespaces}

\bibitem{tamm:1924:jrpc}
\selectlanguageifdefined{russian}
\BibEmph{Тамм~И.~Е.} {Электродинамика
  анизотропной среды в специальной теории
  относительности}~// \BibEmph{Журнал Русского
  физико-химического общества. Часть
  физическая}. \BibDash
\newblock 1924. \BibDash
\newblock \CYRT.~56, {\cyr\textnumero} 2-3. \BibDash
\newblock {\cyr\CYRS.}~248--262.

\bibitem{tamm:1925:jrpc}
\selectlanguageifdefined{russian}
\BibEmph{Тамм~И.~Е.} {Кристаллооптика теории
  относительности в связи с геометрией
  биквадратичной формы}~// \BibEmph{Журнал
  Русского физико-химического общества.
  Часть физическая}. \BibDash
\newblock 1925. \BibDash
\newblock \CYRT.~57, {\cyr\textnumero} 3-4. \BibDash
\newblock {\cyr\CYRS.}~209--240.

\bibitem{tamm:1925:mathann}
\selectlanguageifdefined{german}
\BibEmph{Tamm~I.~E., Mandelstam~L.~I.} {Elektrodynamik der anisotropen Medien
  in der speziellen Relativitatstheorie}~// \BibEmph{Mathematische Annalen}.
  \BibDash
\newblock 1925. \BibDash
\newblock {Bd.}~95, {H.}~1. \BibDash
\newblock S.~154--160.

\bibitem{Plebanski1960}
\selectlanguageifdefined{english}
\BibEmph{Plebanski~J.} {Electromagnetic waves in gravitational fields}~//
  \href{http://dx.doi.org/10.1103/PhysRev.118.1396}{\BibEmph{Physical Review}}.
  \BibDash
\newblock 1960. \BibDash
\newblock Vol. 118, no.~5. \BibDash
\newblock P.~1396--1408. \BibDash
\newblock URL: \BibUrl{http://prola.aps.org/abstract/PR/v118/i5/p1396_1}.

\bibitem{Felice1971}
\selectlanguageifdefined{english}
\BibEmph{Felice~F.} {On the gravitational field acting as an optical medium}~//
  \href{http://dx.doi.org/10.1007/BF00758153}{\BibEmph{General Relativity and
  Gravitation}}. \BibDash
\newblock 1971. \BibDash
\newblock Vol.~2, no.~4. \BibDash
\newblock P.~347--357. \BibDash
\newblock URL: \BibUrl{http://link.springer.com/10.1007/BF00758153}.

\bibitem{Leonhardt2008}
\selectlanguageifdefined{english}
\BibEmph{Leonhardt~U., Philbin~T.~G., Haugh~N.} {General Relativity in
  Electrical Engineering}. \BibDash
\newblock 2008. \BibDash
\newblock P.~1--19. \BibDash
\newblock arXiv~: 0607418v2.

\bibitem{leonhardt:2009:light}
\selectlanguageifdefined{english}
\BibEmph{Leonhardt~U., Philbin~T.~G.}
  \href{http://dx.doi.org/10.1016/S0079-6638(08)00202-3}{{Transformation optics
  and the geometry of light}}~// Progress in Optics. \BibDash
\newblock 2009. \BibDash
\newblock Vol.~53. \BibDash
\newblock P.~69--152. \BibDash
\newblock arXiv~:
  \href{http://www.sciencedirect.com/science/article/pii/S0079663808002023}{0805.4778v2}.

\bibitem{Thompson2011}
\selectlanguageifdefined{english}
\BibEmph{Thompson~R.~T., Cummer~S.~A., Frauendiener~J.} {A completely covariant
  approach to transformation optics}~//
  \href{http://dx.doi.org/10.1088/2040-8978/13/2/024008}{\BibEmph{Journal of
  Optics}}. \BibDash
\newblock 2011. \BibDash
\newblock Vol.~13, no.~2. \BibDash
\newblock P.~024008.

\bibitem{kulyabov:2011:vestnik:curve-maxwell}
\selectlanguageifdefined{russian}
\BibEmph{Кулябов~Д.~С., Немчанинова~Н.~А.}
  {Уравнения Максвелла в криволинейных
  координатах}~// \BibEmph{Вестник РУДН. Серия
  «Математика. Информатика. Физика»}. \BibDash
\newblock 2011. \BibDash
\newblock {\cyr\textnumero}~2. \BibDash
\newblock {\cyr\CYRS.}~172--179.

\bibitem{kulyabov:2012:vestnik:2012-1}
\selectlanguageifdefined{english}
\BibEmph{Kulyabov~D.~S., Korolkova~A.~V., Korolkov~V.~I.} {Maxwell’s
  Equations in Arbitrary Coordinate System}~// \BibEmph{Bulletin of Peoples’
  Friendship University of Russia. Series ``Mathematics. Information Sciences.
  Physics''}. \BibDash
\newblock 2012. \BibDash
\newblock no.~1. \BibDash
\newblock P.~96--106. \BibDash
\newblock arXiv~: \href{http://arxiv.org/abs/1211.6590}{1211.6590}.

\bibitem{penrose-rindler-1987}
\selectlanguageifdefined{russian}
\BibEmph{Пенроуз~Р., Риндлер~В.} {Спиноры и
  пространство-время. Два-спинорное
  исчисление и релятивистские поля}. \BibDash
\newblock М.~: Мир, 1987. \BibDash
\newblock \CYRT.~1. \BibDash
\newblock 528~{\cyr\cyrs.}

\bibitem{sivukhin:1979:ufn}
\selectlanguageifdefined{russian}
\BibEmph{Сивухин~Д.~В.} {О Международной системе
  физических величин}~//
  \href{http://dx.doi.org/10.3367/UFNr.0129.197910h.0335}{\BibEmph{Успехи
  физических наук}}. \BibDash
\newblock 1979. \BibDash
\newblock \CYRT. 129, {\cyr\textnumero}~10. \BibDash
\newblock {\cyr\CYRS.}~335--338. \BibDash
\newblock URL: \BibUrl{http://ufn.ru/ru/articles/1979/10/h/}.

\bibitem{kulyabov:2013:springer:cadabra}
\selectlanguageifdefined{english}
\BibEmph{Korol’kova~A.~V., Kulyabov~D.~S., Sevast’yanov~L.~A.} {Tensor
  computations in computer algebra systems}~//
  \href{http://dx.doi.org/10.1134/S0361768813030031}{\BibEmph{Programming and
  Computer Software}}. \BibDash
\newblock 2013. \BibDash
\newblock Vol.~39, no.~3. \BibDash
\newblock P.~135--142. \BibDash
\newblock URL: \BibUrl{http://link.springer.com/10.1134/S0361768813030031}.

\bibitem{kulyabov:2013:conf:maxwell}
\selectlanguageifdefined{english}
\BibEmph{Kulyabov~D.~S.} {Geometrization of Electromagnetic Waves}~//
  Mathematical Modeling and Computational Physics. \BibDash
\newblock Dubna~: JINR, 2013. \BibDash
\newblock P.~120. \BibDash
\newblock URL: \BibUrl{http://mmcp2013.jinr.ru}.

\bibitem{kulyabov:2013:tver:maxwell}
\selectlanguageifdefined{russian}
\BibEmph{Кулябов~Д.~С., Королькова~А.~В.}
  {Уравнения Максвелла в произвольной
  системе координат}~// \BibEmph{Вестник
  Тверского государственного университета.
  Серия: Прикладная математика}. \BibDash
\newblock 2013. \BibDash
\newblock {\cyr\textnumero} 1 (28). \BibDash
\newblock {\cyr\CYRS.}~29--44.

\bibitem{minkowski:1908}
\selectlanguageifdefined{german}
\BibEmph{Minkowski~H.} {Die Grundlagen f\"{u}r die electromagnetischen
  Vorg\"{a}nge in bewegten K\"{o}rpern}~// \BibEmph{Nachrichten von der
  Gesellschaft der Wissenschaften zu G\"{o}ttingen, Mathematisch-Physikalische
  Klasse}. \BibDash
\newblock 1908. \BibDash
\newblock {H.}~68. \BibDash
\newblock S.~53--111.

\bibitem{stratton:1948}
\selectlanguageifdefined{russian}
\BibEmph{Стрэттон~Д.~А.} Теория
  электромагнетизма. \BibDash
\newblock М.-Л.: ГИТТЛ, 1948.

\bibitem{terletskiy-rybakov-1990}
\selectlanguageifdefined{russian}
\BibEmph{Терлецкий~Я.~П., Рыбаков~Ю.~П.}
  Электродинамика. \BibDash
\newblock 2-е, перераб. {\cyr\cyri\cyrz\cyrd.} \BibDash
\newblock М.~: Высшая школа, 1990. \BibDash
\newblock 352~{\cyr\cyrs.}

\bibitem{sommerfeld1958}
\selectlanguageifdefined{russian}
\BibEmph{Зоммерфельд~А.} Электродинамика. \BibDash
\newblock М.~: Издательство иностранной
  литературы, 1958.

\end{thebibliography}

\RequirePackage{fix-cm}

\documentclass[%
twocolumn, %
nofootinbib,
]{revtex4-1}
\usepackage{cmap}

\usepackage{ucs}
\usepackage[utf8x]{inputenc}
\usepackage[T1,T2A]{fontenc}
\usepackage[german,russian,english]{babel}

\usepackage[sort&compress]{natbib}
\usepackage{amsmath}
\usepackage{amssymb}

\usepackage{mathtools}
\mathtoolsset{
showonlyrefs,
mathic = true
}

\allowdisplaybreaks

\usepackage{hyperref}
\hypersetup{backref,
 colorlinks=false}
\hypersetup{pdfborder=0 0 0}

\usepackage{breakurl}

\usepackage{microtype}
\UseMicrotypeSet[protrusion]{alltext}

\usepackage{fixltx2e}

\usepackage{nicefrac}

\usepackage{physymb}

\makeatletter
\def\ps@pprintTitle{%
     \let\@oddhead\@empty
     \let\@evenhead\@empty
     \let\@oddfoot\@empty
     \let\@evenfoot\@oddfoot}
\makeatother

\newcommand{\crd}[1]{\underline{\vphantom{j}{#1}}}

\begin{document}

\title{The simplest geometrization of Maxwell's equations}

\author{D. S. Kulyabov}
\email{yamadharma@gmail.com}
\author{A. V. Korolkova}
\email{avkorolkova@gmail.com} 
\author{L. A. Sevastyanov}
\email{leonid.sevast@gmail.com}

\affiliation{Telecommunication System Department\\
Peoples' Friendship University of Russia\\
Miklukho-Maklaya str., 6, Moscow, 117198, Russia}


\thanks{Sources:
  \url{https://bitbucket.org/yamadharma/articles-2013-geom-maxwell}}

\begin{abstract}







  For research in the field of transformation optics and for the
  calculation of optically inhomogeneous lenses the method of
  geometrization of the Maxwell equations seems to be perspective. The
  basic idea is to transform the coefficients of material equations,
  namely the dielectric permittivity and magnetic permeability in the
  effective geometry of space-time (besides the vacuum Maxwell
  equations). This allows us to solve the direct and inverse problems,
  that is, to find the permittivity and magnetic permeability for a
  given effective geometry (paths of rays), as well as finding an
  effective geometry on the dielectric permittivity and magnetic
  permeability. The most popular naive geometrization was proposed by
  J. Plebanski. Under certain limitations it is quite good for solving
  relevant problems. It should be noted that in his paper only the
  resulting formulas and exclusively for Cartesian coordinate systems
  are given. In our work we conducted a detailed derivation of
  formulas for the naive geometrization of Maxwell's equations, and
  these formulas are written for an arbitrary curvilinear coordinate
  system. This work is a step toward building a complete covariant
  geometrization of the macroscopic Maxwell’s equations.

\end{abstract}

  \keywords{Maxwell's equations; constitutive equations;
     Maxwell's equations geometrization; Riemann geometry;
     curvilinear coordinates;
     Plebanski's geometrization}

\maketitle

\section{Introduction}
\label{sec:intro}

Differential geometry was an important language of physics
XX-th century. Basic elements of it where developed within the general 
relativity theory. There is a desire to use its power in other
areas of physics, in particular in the optics.

The first attempts to apply the methods of differential geometry in
electrodynamics should be attributed to publications
I.~E.~Tamm~\cite{tamm:1924:jrpc::en, tamm:1925:jrpc::en,
   tamm:1925:mathann}. In 1960 E.~Plebansky proposed method of
geometrization of the constitutive equations of the electromagnetic field, which became
classic~\cite{Plebanski1960, Felice1971, Leonhardt2008,
   leonhardt:2009:light}. All subsequent works, either used it
or tried to correct a little, without changing
ideology~\cite{Thompson2011}. Unfortunately Plebansky~\cite{Plebanski1960} gives no
deriving formulas. Ideology is not expressed 
explicitly too.

For applying and deepening geometrization of material
equations the authors have restored the ideology and specific
Plebanski's calculations.

In paragraph~\ref{sec:notation} we prosecuted provides basic notation and conventions
used in the article. In paragraph~\ref{sec:maxwell_curv} are the main
relations for the Maxwell's equations in curvilinear coordinates (for
more detailed discussion the reader can be refer to other 
authors articles~\cite{kulyabov:2011:vestnik:curve-maxwell::en,
  kulyabov:2012:vestnik:2012-1}). In paragraph~\ref{sec:formal_geometr} are presented actual
calculations on Plebanski geometrization.

\section{Notations and conventions}
\label{sec:notation}

  \begin{enumerate}

  \item We will use the notation of abstract
    indices~\cite{penrose-rindler-1987::en}. In this notation tensor
    as a 
    complete object is denoted merely by an index (eg, $x^{i}$). Its
    components are
    designated by underlined indices (e.g., $x^{\crd{i}}$).

  \item We will adhere to the following agreements . Greek indices
    ($\alpha$, $\beta$) will refer to the four-dimensional space , in
    component form it looks like: $\crd{\alpha} = \overline{0,3}$.  Latin
    indices from the middle of the alphabet ($i$, $j$, $k$) will refer
    to the three-dimensional space , in the component form it looks like:
    $\crd {i} = \overline{1,3} $.

  \item The comma in the index denotes partial derivative with respect to
    corresponding coordinate ($f_{, i} := \partial_{i}f$); semicolon
    denotes covariant derivative ($f_{;i} := \nabla_{i} f$).

  \item To write the equations of electrodynamics in the article is used
    CGS symmetrical system~\cite{sivukhin:1979:ufn::en}.

  \item Antisymmetrization is denoted by straight brackets.

  \end{enumerate}

\section{Maxwell's equations in curvilinear coordinates}
\label{sec:maxwell_curv}

Here are the basics of Maxwell's equations in curvilinear
coordinates. A more detailed description is given in
articles~\cite{kulyabov:2012:vestnik:2012-1, kulyabov:2013:springer:cadabra,
  kulyabov:2013:conf:maxwell,
  kulyabov:2011:vestnik:curve-maxwell::en}.

Let's write the Maxwell equations with the help of electromagnetic
field tensors $F_{\alpha\beta}$ and
$G_{{\alpha}{\beta}}$~\cite{minkowski:1908,stratton:1948::en}:
\begin{gather}
\nabla_{{\alpha}} F_{{\beta}{\gamma}}+ \nabla_{{\beta}}
F_{{\gamma}{\alpha}}+\nabla_{{\gamma}} F_{{\alpha}{\beta}} = 
F_{[\alpha \beta ; \gamma]} = 0,
\label{eq:m:tensor:2}
\\
\nabla_{{\alpha}} G^{{\alpha}{\beta}}=\frac{4 \pi}{c}j^{{\beta}},
\label{eq:m:tensor}
\end{gather}
where the tensors $F_{\alpha \beta}$ and $G^{\alpha \beta}$ have the following
components
\begin{gather}
F_{\crd{\alpha}\crd{\beta}}=
\begin{pmatrix} 
0 & {E}_1 & {E}_2 & {E}_3 \\ 
-{E}_1 & 0 & -{B}^3 & {B}^2 \\
-{E}_2 & {B}^3 & 0 & -{B}^1 \\ 
-{E}_3 & -{B}^2 & {B}^1 & 0 
\end{pmatrix},
\label{eq:f_ab}
\\
F^{\crd{\alpha}\crd{\beta}}=
\begin{pmatrix} 
0 & -{E}^1 & -{E}^2 & -{E}^3 \\ 
{E}^1 & 0 & -{B}_3 & {B}_2 \\
{E}^2 & {B}_3 & 0 & -{B}_1 \\ 
{E}^3 & -{B}_2 & {B}_1 & 0 
\end{pmatrix},
\label{eq:f^ab}
\\
G^{\crd{\alpha}\crd{\beta}}=
\begin{pmatrix} 
0 & -{D}^1 & -{D}^2 & -{D}^3 \\ 
{D}^1 & 0 & -{H}_3 & {H}_2 \\
{D}^2 & {H}_3 & 0 & -{H}_1 \\ 
{D}^3 & -{H}_2 & {H}_1 & 0 
\end{pmatrix},
\label{eq:g^ab}
\\
G_{\crd{\alpha}\crd{\beta}}=
\begin{pmatrix} 
0 & {D}_1 & {D}_2 & {D}_3 \\ 
-{D}_1 & 0 & -{H}^3 & {H}^2 \\
-{D}_2 & {H}^3 & 0 & -{H}^1 \\ 
-{D}_3 & -{H}^2 & {H}^1 & 0 
\end{pmatrix}.
\label{eq:g_ab}
\end{gather}
  Where $E_{\crd{i}}$, $H_{\crd{i}}$ are components of electric and magnetic
  fields intensity vectors; $D^{\crd{i}}$, $B^{\crd{i}}$ are
  components of vectors of electric and magnetic induction.

It is also useful to introduce the tensor $\prescript{*}{}{F}^{\alpha\beta}$ dual conjugated to
${F}_{\alpha\beta}$
\begin{equation}
  \label{eq:tilde_f}
  \prescript{*}{}{F}^{\alpha\beta} = \frac{1}{2} e^{\alpha\beta\gamma\delta} F_{\gamma\delta},
\end{equation}
where $e^{\alpha\beta\gamma\delta}$ is the alternating tensor,
$\varepsilon^{\alpha\beta\gamma\delta}$ is the Levi-Civita symbol:
\begin{equation}
  \label{eq:e4}
  e_{\crd{\alpha}\crd{\beta}\crd{\gamma}\crd{\delta}} = \sqrt{-g}\varepsilon_{\crd{\alpha}\crd{\beta}\crd{\gamma}\crd{\delta}},\quad
  e^{\crd{\alpha}\crd{\beta}\crd{\gamma}\crd{\delta}} = -\frac{1}{\sqrt{-g}}\varepsilon^{\crd{\alpha}\crd{\beta}\crd{\gamma}\crd{\delta}}.
\end{equation}

Similarly, we can write
\begin{equation}
  \label{eq:tilde_other}
    \prescript{*}{}{G}_{\alpha\beta} = \frac{1}{2}
    e_{\alpha\beta\gamma\delta} G^{\gamma\delta}, 
\end{equation}
The components of these dual tensors are given (by the expressions) as follows
\begin{gather}
\prescript{*}{}{F}^{\crd{\alpha}\crd{\beta}}=
\frac{1}{\sqrt{-g}}
\begin{pmatrix} 
0 & -{B}^1 & -{B}^2 & -{B}^3 \\ 
{B}^1 & 0 & {E}_3 & -{E}_2 \\
{B}^2 & -{E}_3 & 0 & {E}_1 \\ 
{B}^3 & {E}_2 & -{E}_1 & 0 
\end{pmatrix},
\label{eq:*f^ab}
\\
\prescript{*}{}{G}_{\crd{\alpha}\crd{\beta}}=
\sqrt{-g}
\begin{pmatrix} 
0 & {H}_1 & {H}_2 & {H}_3 \\ 
-{H}_1 & 0 & {D}^3 & -{D}^2 \\
-{H}_2 & -{D}^3 & 0 & {D}^1 \\ 
-{H}_3 & {D}^2 & -{D}^1 & 0 
\end{pmatrix}.
\label{eq:*g_ab}
\end{gather}

With dual tensor~\eqref{eq:tilde_f}
the equation \eqref{eq:m:tensor:2} may be rewritten in a simpler form:
\begin{equation}
  \label{eq:m:tensor:2:dual}
  \nabla_{\alpha} \prescript{*}{}{F}^{\alpha\beta} = 0.
\end{equation}

Next, we will explain why, for the purposes of implementation of the
Plebanski's program, we prefer to write equations in the
form~\eqref{eq:m:tensor:2} rather the in simpler
form~\eqref{eq:m:tensor:2:dual}.

\subsection{Maxwell's equations in a medium}

In the presence of the medium a group of equations
Maxwell containing bound charges is changed, namely
equation~\eqref{eq:m:tensor}.
Amongst the Maxwell's equations~\eqref{eq:m:tensor:2}
and~\eqref{eq:m:tensor} must be added the constitutive relations between
tensors $G^{\alpha \beta}$ and $F^{\alpha \beta}$. When introduced
additional assumptions about the linearity of the environment and immobility
substances they may be written in three-dimensional form as follows:
\begin{equation}
  \label{eq:constraint}
  D^{i} = \varepsilon^{i j} E_{j}, \qquad B^{i} = \mu^{i j} H_{j},
\end{equation}
where $\varepsilon^{ij}$ and $\mu^{ij}$ are the permittivity and
permeability tensors. In four-dimensional form relation~\eqref{eq:constraint}
takes the following form:
\begin{equation}
  \label{eq:constraint:4}
  G^{\alpha \beta} = \lambda^{\alpha \beta}_{\gamma \delta} F^{\gamma
    \delta}, 
  \quad 
  \lambda^{\alpha \beta}_{\gamma \delta} = \lambda^{[\alpha \beta]}_{[\gamma \delta]},
\end{equation}
here tensor $\lambda^{\alpha \beta}_{\gamma \delta}$ is containing
information on the permittivity and permeability, as well as
electro-magnetic coupling~\cite{tamm:1924:jrpc::en}.

Given the structure tensor $F^{\alpha \beta}$~\eqref{eq:f^ab} and
$G^{\alpha \beta}$~\eqref{eq:g^ab}, and the
constitutive relations~\eqref{eq:constraint}, we write
\begin{equation}
  \label{eq:f-g_lambda}
  \begin{gathered}
    F^{\crd{0}\crd{i}} = - E^{\crd{i}}, \quad
    G^{\crd{0}\crd{i}} = - D^{\crd{i}},
    \\
    G^{\crd{i}\crd{j}} = - \varepsilon^{\crd{i} \crd{j} \crd{k}}
    H_{\crd{k}}, \quad
    B_{\crd{i}} = - \varepsilon_{\crd{i} \crd{j} \crd{k}} F^{\crd{j}
      \crd{k}}.
  \end{gathered}
\end{equation}
Or in another form:
\begin{equation}
  \label{eq:f-g_lambda_2}
  G^{\crd{0}\crd{i}} = \varepsilon^{\crd{i}}_{\crd{j}} F^{0 \crd{j}},
  \quad
  G^{\crd{i}\crd{j}} = \varepsilon^{\crd{i} \crd{j} \crd{k}}
  \varepsilon_{\crd{l} \crd{m} \crd{n}}
  (\mu^{-1})^{\crd{l}}_{\crd{k}} F^{\crd{m} \crd{n}}/2.
\end{equation}

From the~\eqref{eq:f-g_lambda_2} we write the structure of tensor
$\lambda^{\alpha \beta}_{\gamma \delta}$:
\begin{equation}
  \label{eq:lambda_struct}
  \lambda^{0 \crd{i}}_{0 \crd{j}} = \varepsilon^{\crd{i}}_{\crd{j}} /2, \quad
  \lambda^{0 \crd{i}}_{\crd{k} \crd{l}} = \lambda_{0 \crd{i}}^{\crd{k}
    \crd{l}} = 0, \quad
  \lambda^{\crd{i} \crd{j}}_{\crd{m} \crd{n}} = 
  \varepsilon^{\crd{i} \crd{j} \crd{k}}
  \varepsilon_{\crd{l} \crd{m} \crd{n}}
  (\mu^{-1})^{\crd{l}}_{\crd{k}}/2.
\end{equation}

Assuming that the vacuum permittivity and permeability are
view:
\begin{equation}
  \label{eq:perv:vac}
  \varepsilon^{i j} := \delta^{i j}, \qquad \mu^{i j} := \delta^{i j},
\end{equation}
we find that the vacuum constitutive relations~\eqref{eq:constraint} and~\eqref{eq:constraint:4} take
view
\begin{equation}
  \label{eq:constraint:vac}
  D^{i} = E^{i}, \qquad B^{i} = H^{i}, 
  \qquad G^{\alpha \beta} = F^{\alpha \beta}.
\end{equation}

\subsubsection{Permeability tensor in an isotropic medium}
\label{sec:perm_isotropic}

In the case of an isotropic medium, the
expression~\eqref{eq:constraint} takes the form:
\begin{equation}
  \label{eq:constraint:isotropic}
  D^{i} = \varepsilon E^{j}, \qquad B^{i} = \mu H^{j},
\end{equation}
where $\varepsilon^{ij}$ and $\mu^{ij}$ are the permittivity and
permeability scalars.

In this case, the permeability tensor in a stationary frame of
reference can be represented as follows~\cite{tamm:1925:mathann,
  tamm:1924:jrpc::en}:
\begin{equation}
  \label{eq:perm_tensor_2val}
  \begin{gathered}
    \lambda_{\alpha \beta \gamma \delta} = \lambda_{\alpha \gamma}
    \lambda_{\beta \delta},
    \\
    \lambda_{\crd{\alpha} \crd{\beta}} = 
    \mathrm{diag}\left(
      \frac{1}{\varepsilon \sqrt{\mu}}, -\sqrt{\mu}, -\sqrt{\mu},
      -\sqrt{\mu}
      \right), 
      \\
    \lambda^{\crd{\alpha} \crd{\beta}} =
    \mathrm{diag}\left(
      \varepsilon \sqrt{\mu}, -\frac{1}{\sqrt{\mu}}, -\frac{1}{\sqrt{\mu}},
      -\frac{1}{\sqrt{\mu}}
      \right).
  \end{gathered}
\end{equation}

\subsubsection{Constitutive relations in moving media}
\label{sec:moving_media}

Minkowski derived equations for connection
isotropic moving media~\cite{minkowski:1908, sommerfeld1964lectures}
(Minkowski's equations for moving media).
Let $u^{\alpha}$ is a environments four-speed. Assuming that the
permittivity and permeability $\varepsilon$ and $\mu$ are
scalars, we can write
\begin{equation}
  \label{eq:moving_isotropic_4d}
  G^{\alpha \beta} u_{\beta} = \varepsilon F^{\alpha \beta} u_{\beta}, 
  \quad
  \prescript{*}{}{F}^{\alpha \beta} u_{\beta} = \mu 
  \prescript{*}{}{G}^{\alpha \beta} u_{\beta}.
\end{equation}

In three-dimensional form of the
equation~\eqref{eq:moving_isotropic_4d} take the following form:
\begin{equation}
  \label{eq:moving_isotropic_3d}
  \begin{gathered}
    \begin{multlined}
      D^{i} = \varepsilon \left( E^{i} + 
        \left[\frac{u_{j}}{c}, B_{k} \right]^{i}
      \right) -
      \left[\frac{u_{j}}{c}, H_{k} \right]^{i} 
      = {} \\ {} =
      \varepsilon E^{i} + (\varepsilon \mu - 1) 
      \left[\frac{u_{j}}{c}, H_{k} \right]^{i}, 
    \end{multlined}
    \\
    \begin{multlined}
      B^{i} = \mu \left( H^{i} - 
        \left[\frac{u_{j}}{c}, D_{k} \right]^{i}
      \right) +
      \left[\frac{u_{j}}{c}, E_{k} \right]^{i} 
      = {} \\ {} = 
      \mu H^{i} - (\varepsilon \mu - 1) 
      \left[ \frac{u_{j}}{c}, E_{k} \right]^{i}.
    \end{multlined}
  \end{gathered}
\end{equation}

Tamm expanded equation~\eqref{eq:moving_isotropic_3d} for the
anisotropic case~\cite{tamm:1924:jrpc::en,
  tamm:1925:mathann}, namely, assuming that the permittivity and
permeability are of the form
\begin{equation}
  \label{eq:perm_anisotropic}
  \varepsilon^{\crd{i}}_{\crd{j}} = 
  \mathrm{diag}(\varepsilon^{1}_{1}, \varepsilon^{2}_{2},
  \varepsilon^{3}_{3}),
  \quad
  \mu^{\crd{i}}_{\crd{j}} = 
  \mathrm{diag}(\mu^{1}_{1}, \mu^{2}_{2},
  \mu^{3}_{3}),
\end{equation}
and velocity vector $u^{i}$ of frame of reference is parallel to one
of the principal axes of anisotropy. Then the Minkowski's equations
for moving media acquire the following form:
\begin{equation}
  \label{eq:moving_anisotropic_3d}
  \begin{gathered}
    D^{i} = \varepsilon^{i}_{l} \left( E^{l} + 
      \left[\frac{u_{j}}{c}, B_{k} \right]^{l}
    \right) - 
    \left[\frac{u_{j}}{c}, H_{k} \right]^{i} 
    , 
    \\
    B^{i} = \mu^{i}_{l} \left( H^{l} - 
      \left[\frac{u_{j}}{c}, D_{k} \right]^{l}
    \right) + 
    \left[\frac{u_{j}}{c}, E_{k} \right]^{i} 
    .
  \end{gathered}
\end{equation}

\section{Formal geometrization of material Maxwell's equations}
\label{sec:formal_geometr}

Plebanski has offered the elementary geometrization of Maxwell's
equations~\cite{Plebanski1960, Felice1971}. However, in the original
article the final formula are immediately given, and the principles and
methods for their preparation remain obscure. The authors tried to
explicitly describe technique that we believe Plebanski used and to perform
calculations in detail.

The basic ideas of Plebanski's geometrization are as follows:
\begin{enumerate}
\item We write the Maxwell's equations in a medium in Minkowski space.
\item We write the vacuum Maxwell's equations in the effective Riemann space.
\item We equate the corresponding terms of both equations.
\end{enumerate}

As a result, we obtain an expression of the permittivity and
permeability in terms of geometric objects.

Before attempting the program of Plebanski let us recall some auxiliary relations.

\subsection{Auxiliary relations}

\subsubsection{Differential Bianchi identity}

Note that the equation~\eqref{eq:m:tensor:2} can be written as

\begin{equation}
\label{eq:m:tensor:2:partial}
\partial_{{\alpha}} F_{{\beta}{\gamma}}+ \partial_{{\beta}}
F_{{\gamma}{\alpha}}+\partial_{{\gamma}} F_{{\alpha}{\beta}} = 
F_{[\alpha \beta , \gamma]} = 0.
\end{equation}

Really:
\begin{multline}
\label{eq:m:tensor:2:detail}
\nabla_{{\alpha}} F_{{\beta}{\gamma}}+ \nabla_{{\beta}}
F_{{\gamma}{\alpha}}+\nabla_{{\gamma}} F_{{\alpha}{\beta}} = {} \\ {} =
\partial_{{\alpha}} F_{{\beta}{\gamma}} -
\Gamma_{\alpha\beta}^{\delta} F_{\delta \gamma} - 
\Gamma_{\alpha \gamma}^{\delta} F_{\beta \delta} +
\partial_{{\beta}} F_{{\gamma}{\alpha}} -
\Gamma_{\beta \gamma}^{\delta} F_{\delta \alpha} -
\Gamma_{\beta \alpha}^{\delta} F_{\gamma \delta}
+ {} \\ {} +
\partial_{{\gamma}} F_{{\alpha}{\beta}} -
\Gamma_{\gamma \alpha}^{\delta} F_{\delta \beta} -
\Gamma_{\gamma \beta}^{\delta} F_{\alpha \delta}.
\end{multline}
Taking into account the anti-symmetry of the tensor $F_{\alpha \beta}$
and the
symmetry in the lower indices of Christoffel symbols $\Gamma_{\alpha
  \beta}^{\delta}$, we obtain~\eqref{eq:m:tensor:2:partial}.

The resulting equation can be written form-invariant by in any
coordinate system. Therefore, we will use the
equation~\eqref{eq:m:tensor:2:partial} instead of the more commonly
used equation~\eqref{eq:m:tensor:2:dual}.

\subsubsection{The metric tensor relations}

In addition, we need simple relations for the metric
tensor.
\begin{equation}
  \label{eq:g_g^:0}
  g_{\alpha \delta}g^{\delta \beta} = \delta_{\alpha}^{\beta}
\end{equation}
The relation~\eqref{eq:g_g^:0} leads to the following special relations:
\begin{gather}
  g_{0 \delta}g^{\delta i} = g_{0 0}g^{0 i} + g_{0 k}g^{k i}  = \delta_{0}^{i} = 0,
  \label{eq:g_g^:2}
  \\
  g_{i \delta}g^{\delta j} = g_{i 0}g^{0 j} + g_{i k}g^{k j}
  = \delta_{i}^{j}.
  \label{eq:g_g^:3}
\end{gather}
Let equation~\eqref{eq:g_g^:2} be rewritten in
form
\begin{gather}
  g^{0 i} = - \frac{1}{g_{0 0}} g_{0 k}g^{k i}.
  \label{eq:g_g^:2a}
\end{gather}
Substituting equation~\eqref{eq:g_g^:2a} in
equation~\eqref{eq:g_g^:3}, we obtain:
\begin{gather}
  \left(
    g_{i k} - \frac{1}{g_{0 0}} g_{0 i} g_{0 k}
  \right) g^{k j}
  = \delta_{i}^{j}.
  \label{eq:g_g^:2b}
\end{gather}

This relation will be used later to simplify the final equations.

\subsection{Geometrization in Cartesian coordinates}

Let us write the Maxwell's equations in a medium in Cartesian
coordinates with the metric tensor $\eta_{\crd{\alpha} \crd{\beta}} =
\mathrm{diag}(1,-1,-1,-1)$:
\begin{equation}
    \label{eq:maxwell:media:decart}
    \begin{gathered}
      \partial_{{\alpha}} F_{{\beta}{\gamma}} + 
      \partial_{{\beta}} F_{{\gamma}{\alpha}} + 
      \partial_{{\gamma}} F_{{\alpha}{\beta}} = 0,
      \\
      \partial_{{\alpha}} G^{{\alpha}{\beta}} = 
      \frac{4 \pi}{c} j^{{\beta}}.
    \end{gathered}
  \end{equation}

Now we write the vacuum Maxwell's equations in effective Riemann space
with the metric tensor $g_{\alpha \beta}$ (their related values mark
on by the tilde):
\begin{equation}
    \label{eq:maxwell:vacuum:riemann}
    \begin{gathered}
      \partial_{{\alpha}} \Tilde{F}_{{\beta}{\gamma}} + 
      \partial_{{\beta}} \Tilde{F}_{{\gamma}{\alpha}} + 
      \partial_{{\gamma}} \Tilde{F}_{{\alpha}{\beta}} = 0,
      \\
      \frac{1}{\sqrt{-g}}
      \partial_{{\alpha}} 
      \left( \sqrt{-g} 
        \Tilde{G}^{{\alpha}{\beta}} 
      \right)
        = 
      \frac{4 \pi}{c} \tilde{j}^{{\beta}}.
    \end{gathered}
  \end{equation}

In a vacuum, the following relation is true
(see~\eqref{eq:constraint:vac}):
\begin{equation}
  \label{eq:f-g:relation}
  \Tilde{F}_{\alpha \beta} = \Tilde{G}_{\alpha \beta}.
\end{equation}

Raising the indices in~\eqref{eq:f-g:relation}, we obtain
\begin{equation}
  \label{eq:f^-g:relation}
  \Tilde{F}^{\alpha \beta} = g^{\alpha \gamma} g^{\beta \delta}
  \Tilde{G}_{\gamma \delta}.
\end{equation}

Comparing term by term~\eqref{eq:maxwell:media:decart}
and~\eqref{eq:maxwell:vacuum:riemann}, and taking into account
the~\eqref{eq:f^-g:relation} we obtain:
\begin{gather}
  \label{eq:decart-riemann:relation}
  F_{\alpha \beta} = \Tilde{F}_{\alpha \beta}, 
  \qquad
  j^{\alpha} = \sqrt{-g} \Tilde{j}^{\alpha},
  \\
  G^{\alpha \beta}
  = \sqrt{-g} g^{\alpha \gamma} g^{\beta \delta} F_{\gamma \delta}.
  \label{eq:decart-riemann:relation:g-f}
\end{gather}

Equations~\eqref{eq:decart-riemann:relation:g-f} are actually
4-dimensional geometrized constitutive relations~\eqref{eq:constraint:4},
we were seeking for. Following the Plebanski's program 
we must obtain the explicit form for the 3-dimensional constitutive
relations~\eqref{eq:constraint}.

\subsubsection{Electric displacement field}

Let us rewrite~\eqref{eq:decart-riemann:relation:g-f} in the form of:
\begin{equation}
  \label{eq:f-durch-g}
  F_{\alpha \beta} = 
  \frac{1}{\sqrt{-g}} g_{\alpha \gamma} g_{\beta \delta} 
  G^{\gamma \delta}
\end{equation}
and look for the value components of $F_{0 \crd{i}}$, taking into
account relations~\eqref{eq:f_ab} and~\eqref{eq:g^ab}:
\begin{multline}
  \label{eq:f_0i}
  F_{0 i} = E_{i} = \frac{1}{\sqrt{-g}} 
  g_{0 \gamma} g_{i \delta} G^{\gamma \delta} 
  = {} \\ {} = 
  \frac{1}{\sqrt{-g}} \left(
  g_{0 j} g_{i 0} G^{j 0} + g_{0 0} g_{i j} G^{0 j}
  \right) +
  \frac{1}{\sqrt{-g}} g_{0 j} g_{i k} G^{j k} 
  = {} \\ {} = 
  \frac{1}{\sqrt{-g}} 
  g_{0 0} \left(
    \frac{1}{g_{0 0}} g_{0 j} g_{i 0} - g_{i j} 
  \right) D^{j} -
  \frac{1}{\sqrt{-g}} g_{0 j} g_{i k} \varepsilon^{j k l} H_{l}.
\end{multline}
For induction components $D^{i}$ we apply the relation~\eqref{eq:g_g^:2b}
and obtain
\begin{equation}
  \label{eq:d^i}
  D^{i} = - \frac{\sqrt{-g}}{g_{0 0}} g^{i j} E_{j} + 
  \frac{1}{g_{0 0}} \varepsilon^{i j k} g_{j 0} H_{k}.
\end{equation}

From~\eqref{eq:d^i} we can formally deduce the expression for the
permittivity:
\begin{equation}
  \label{eq:e_ij} 
  \varepsilon^{i j} = - \frac{\sqrt{-g}}{g_{0 0}} g^{i j}.
\end{equation}
In this sense the second term in~\eqref{eq:d^i} needs further
clarification.

\subsubsection{Magnetic induction}

To obtain expressions for the magnetic induction we will
use tensors~\eqref{eq:*f^ab} and~\eqref{eq:*g_ab}.
Moving down the indices of $\prescript{*}{}{G}^{\alpha \beta}$
in terms of~\eqref{eq:decart-riemann:relation:g-f} and applying
relations \eqref{eq:tilde_f}, \eqref{eq:tilde_other} we obtain
\begin{equation}
  \label{eq:*g-durch-*f}
  \prescript{*}{}{G}_{\alpha \beta} = 
  \sqrt{-g} g_{\alpha \gamma} g_{\beta \delta} 
  \prescript{*}{}{F}^{\gamma \delta}.
\end{equation}

We will seek for the value components of $\prescript{*}{}{G}_{0
  \crd{i}}$:
\begin{multline}
  \label{eq:*g_0i}
  \prescript{*}{}{G}_{0 i} = \sqrt{-g} H_{i} = 
  \sqrt{-g} g_{0 \gamma} g_{i \delta} 
  \prescript{*}{}{F}^{\gamma \delta} 
  = {} \\ {} = 
  \sqrt{-g} \left(
  g_{0 j} g_{i 0} \prescript{*}{}{F}^{j 0} + 
  g_{0 0} g_{i j} \prescript{*}{}{F}^{0 j}
  \right) +
  \sqrt{-g} g_{0 j} g_{i k} 
  \prescript{*}{}{F}^{j k} 
  = {} \\ {} = 
  \sqrt{-g}
  g_{0 0} \left(
    \frac{1}{g_{0 0}} g_{0 j} g_{i 0} - g_{i j} 
  \right) \frac{1}{\sqrt{-g}} B^{j} 
  - {} \\ {} -
  \sqrt{-g} g_{0 j} g_{i k} \varepsilon^{j k l} 
  \frac{1}{\sqrt{-g}} E_{l}.
\end{multline}

Applying the relation~\eqref{eq:g_g^:2b} we obtain for $B^{i}$ the
following expression:
\begin{equation}
  \label{eq:b^i}
  B^{i} = - \frac{\sqrt{-g}}{g_{0 0}} g^{i j} H_{j} - 
  \frac{1}{g_{0 0}} \varepsilon^{i j k} g_{j 0} E_{k}.
\end{equation}

From~\eqref{eq:b^i} we can formally write the expression for
permeability:
\begin{equation}
  \label{eq:mu_ij} 
  \mu^{i j} = - \frac{\sqrt{-g}}{g_{0 0}} g^{i j}.
\end{equation}

Thus geometrized constitutive relations in Cartesian
coordinates are as follows:
\begin{equation}
  \label{eq:geom-maxwell:decart}
  \begin{gathered}
    D^{i} = \varepsilon^{i j} E_{j} + 
    \varepsilon^{i j k} w_{j} H_{k}, \\
    B^{i} = \mu^{i j} H_{j} - 
    \varepsilon^{i j k} w_{j} E_{k}, \\
    \varepsilon^{i j} = - \frac{\sqrt{-g}}{g_{0 0}} g^{i j}, \quad
    \mu^{i j} = - \frac{\sqrt{-g}}{g_{0 0}} g^{i j}, \quad 
    w_{i} = \frac{g_{i 0}}{g_{0 0}}.
  \end{gathered}
\end{equation}

These equations were given in the original
paper~\cite{Plebanski1960}. For now, we can assume that we have performed
our task.

\subsubsection{Electro-magnetic interaction term interpretation}

In the equation~\eqref{eq:geom-maxwell:decart} the term of
electro-magnetic interactions does
receive no interpretation in original Plebanski's article. However
Leonhard proposed to interpret it as a speed
of geometrized frame of reference~\cite {Leonhardt2008}. Indeed, on
the basis of~\eqref{eq:moving_isotropic_3d}
equation~\eqref{eq:geom-maxwell:decart} can be rewritten as:
\begin{equation}
  \label{eq:geom-maxwell:decart:moving}
  \begin{gathered}
    D^{i} = \varepsilon^{i j} E_{j} + 
    \left[ \frac{u_{j}}{c}, H_{k} \right]^{i}, \\
    B^{i} = \mu^{i j} H_{j} - 
    \left[ \frac{u_{j}}{c}, E_{k} \right]^{i}, \\
    \varepsilon^{i j} = - \frac{\sqrt{-g}}{g_{0 0}} g^{i j}, \quad
    \mu^{i j} = - \frac{\sqrt{-g}}{g_{0 0}} g^{i j}, \quad 
    u_{i} = \frac{g_{i 0}}{g_{0 0}} \frac{c \sqrt{g^{(3)}}}{n^2 - 1},
  \end{gathered}
\end{equation}
where $u^{i}$ is three-dimensional velocity of frame of reference,
$u^{i}$ is determinant of the spatial part of
metric tensor $g_{\alpha \beta}$,
$n = \sqrt{\varepsilon \mu}$ is refractive index.

\subsection{Geometrization in curvilinear coordinates}

We now extend the scope of the formulas obtained to the Maxwell's
equations in arbitrary curvilinear
coordinates. Suppose that this space is defined by the metric
tensor $\gamma_{\alpha \beta}$. Then the
system~\eqref{eq:maxwell:media:decart} takes the following form:
\begin{equation}
    \label{eq:maxwell:media:curve}
    \begin{gathered}
      \partial_{{\alpha}} F_{{\beta}{\gamma}} + 
      \partial_{{\beta}} F_{{\gamma}{\alpha}} + 
      \partial_{{\gamma}} F_{{\alpha}{\beta}} = 0,
      \\
      \frac{1}{\sqrt{-\gamma}} 
      \partial_{{\alpha}} \left(
      \sqrt{-\gamma} G^{{\alpha}{\beta}} 
      \right)
      = 
      \frac{4 \pi}{c} j^{{\beta}}.
    \end{gathered}
  \end{equation}

Further, repeating the steps for the effective Riemann
space~\eqref{eq:maxwell:vacuum:riemann},
\eqref{eq:f-g:relation}, \eqref{eq:f^-g:relation}, we obtain the
analogues to~\eqref{eq:decart-riemann:relation}
and~\eqref{eq:decart-riemann:relation:g-f} as follows
\begin{gather}
  \label{eq:curv-riemann:relation}
  F_{\alpha \beta} = \Tilde{F}_{\alpha \beta}, 
  \qquad 
  j^{\alpha} = \frac{\sqrt{-g}}{\sqrt{-\gamma}} \Tilde{j}^{\alpha},
  \\
  G^{\alpha \beta}
  = \frac{\sqrt{-g}}{\sqrt{-\gamma}} g^{\alpha \gamma} g^{\beta \delta} F_{\gamma \delta}.
  \label{eq:curv-riemann:relation:g-f}
\end{gather}

\subsubsection{Electric displacement field}

We write the expression~\eqref{eq:curv-riemann:relation:g-f} as:
\begin{equation}
  \label{eq:f-durch-g:curv}
  F_{\alpha \beta} = 
  \frac{\sqrt{-\gamma}}{\sqrt{-g}} g_{\alpha \gamma} g_{\beta \delta} 
  G^{\gamma \delta}.
\end{equation}

Arguing similarly to~\eqref{eq:f_0i}, we obtain for the components of the
electric displacement field $D^{i}$ the relation:
\begin{equation}
  \label{eq:d^i:curv}
  D^{i} = - \frac{\sqrt{-g}}{\sqrt{-\gamma}}\frac{1}{g_{0 0}} g^{i j} E_{j} + 
  \frac{1}{g_{0 0}} \varepsilon^{i j k} g_{j 0} H_{k}, 
\end{equation}
and the expression for the permittivity takes the form:
\begin{equation}
  \label{eq:e_ij:curv} 
  \varepsilon^{i j} = - \frac{\sqrt{-g}}{\sqrt{-\gamma}} \frac{1}{g_{0 0}} g^{i j}.
\end{equation}

\subsubsection{Magnetic induction}

Let's rewrite~\eqref{eq:*g-durch-*f}
taking into account~\eqref{eq:curv-riemann:relation:g-f}
\begin{equation}
  \label{eq:*g-durch-*f:curv}
  \prescript{*}{}{G}_{\alpha \beta} = 
  \frac{\sqrt{-g}}{\sqrt{\gamma}} g_{\alpha \gamma} g_{\beta \delta} 
  \prescript{*}{}{F}^{\gamma \delta}.
\end{equation}

By analogy with~\eqref{eq:*g_0i}, \eqref{eq:b^i} and~\eqref{eq:mu_ij}
we obtain relations for the $B^{i}$:
\begin{equation}
  \label{eq:b^i:curv}
  B^{i} = - \frac{\sqrt{-g}}{\sqrt{\gamma}}\frac{1}{g_{0 0}} g^{i j} H_{j} - 
  \frac{1}{g_{0 0}} \varepsilon^{i j k} g_{j 0} E_{k},
\end{equation}
and permeability takes the form
\begin{equation}
  \label{eq:mu_ij:curv} 
  \mu^{i j} = - \frac{\sqrt{-g}}{\sqrt{-\gamma}}\frac{1}{g_{0 0}} g^{i j}.
\end{equation}

Thus geometrized constitutive relations in curvilinear
coordinates with the metric tensor $\gamma_{\alpha \beta}$ are of
the following form:
\begin{equation}
  \label{eq:geom-maxwell:curv}
  \begin{gathered}
    D^{i} = \varepsilon^{i j} E_{j} + 
    \varepsilon^{i j k} w_{j} H_{k}, \\
    B^{i} = \mu^{i j} H_{j} - 
    \varepsilon^{i j k} w_{j} E_{k}, \\
    \varepsilon^{i j} = - \frac{\sqrt{-g}}{\sqrt{-\gamma}} \frac{1}{g_{0 0}} g^{i j}, \qquad
    \mu^{i j} = - \frac{\sqrt{-g}}{\sqrt{-\gamma}}\frac{1}{g_{0 0}} g^{i j}, \qquad 
    w_{i} = \frac{g_{i 0}}{g_{0 0}}.
  \end{gathered}
\end{equation}

\section{Conclusions}
\label{sec:concl}

The authors have restored  the program and calculations of Plebanski.
This approach to geometrization appears inconclusive. 
Namely, the very method of obtaining relations for geometrization
(eq.~\eqref{eq:f^-g:relation}) implicitly assumes an
isotropic medium~\eqref{eq:perm_tensor_2val}.
But this does
not prevent on to
use this method for the calculations in the transformational
optics.

\bibliographystyle{abbrvnat}

\bibliography{bib/geom-maxwell/geom-maxwell,bib/geom-maxwell/maxwell-curvecoord,bib/geom-maxwell/maxwell-curvecoord-other}


\end{document}